The unique colors of the cashmere sweaters your firm makes result from heating undyed yarn in a kettle with a dye liquor. The pH (acidity) of the liquor is critical for regulating dye uptake and hence the final color. There are 7 kettles, all of which receive dye liquor from a common source. Past data show that pH varies according to a Normal distribution with μ = 4.32 and σ = 0.118. You use statistical process control to check the stability of the process. Twice each day, the pH of the liquor in each kettle is measured, giving a sample of size 7. The mean pH x is compared with "control limits" given by the 99.7 part of the 68−95−99.7 rule for normal distributions, namely μx ± 3σx. What are the numerical values of these control limits for x? (Round your answers to three decimal places.)

smaller value

larger value
Answer #1

smaller value = μx - 3σx =4.32-3*0.118/sqrt(7)=4.186

larger value = μx + 3σx =4.32+3*0.118/sqrt(7)=4.454